HUBBLE SPACE TELESCOPE DAILY REPORT #2253 PERIOD COVERED: 0000Z (UTC) 11/17/98 - 0000Z (UTC) 11/18/98 Daily Status Report as of 322/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Eleven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.3 Completed STIS/CCD/MA1/MA2 7399 (Ultraviolet Spectroscopy of the Supersoft X-ray Source RX J0439.8-6809) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the super soft x-ray source J0439.8-6809. Observationally, supersoft X-ray sources are classified as near-Eddington stellar objects with almost all emission at energies < 0.5 keV. Only 13 supersoft X-ray sources have been optically identified, and of these 11 turn out to be binaries, probably with a shell-burning accreting white dwarf. A wavelength calibration observations was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Three Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.5 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed WF/PC-2 6347 (A Search for Jets in Planetary Nebulae) The WF/PC-2 was used to make observations of the planetary nebulae (class PNe) NGC-6881 in order to search for the presence of possible narrow jets. The current model for PNe formation indicates that jets are an unexpected phenomena. Yet there is a growing body of circumstantial evidence that suggests narrow jets are common feature in PNe. The observations were completed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 6569 (The Physics of Nearby Blue Irregular Starbursts) The WF/PC-2 was used to make observations of nearby dwarf galaxy UGC-685 in order to study the physics of its blue irregular starburst regions by obtaining ultraviolet, violet, red, and infrared and narrow band H Alpha and O III images. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Five Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.10 Completed NIC/2/3 8082 (NICMOS Pure Parallel Fall 98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel observations while another instrument is prime. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2 6554 (Extragalactic Globular Cluster Systems) The WF/PC-2 was used to make observations of a extragalactic globular cluster system which is located in galaxy NGC-7562. This is part of a survey which will cover a range in galaxy luminosity, morphological type, environment, galactic cluster specific frequency and Galactic cluster mean metallicity. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Six Sets of STIS/CCD 8086 (Leonid STIS proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of quasar PKS2200-23 during the Leonid meteor shower (HST Leonid-avoidance pointing) to identify high redshift galaxies associated with Lyman-alpha and metal absorption systems in the quasar spectrum. Our key aim is to achieve the direct optical identification of high redshift galaxies which produce absorption against the background QSO. The observations were completed as scheduled, and no anomalies were noted. 1.13 Completed Six Sets of WF/PC-2 8087 (Leonid WF/PC-2 proposal) The WF/PC-2 was used to make observations during the Leonid meteor shower. This is a special pure-parallel service observation proposal which will execute during the HST Leonid-avoidance pointing. A deep WF/PC-2 image was taken over 6 orbits {in parallel with a STIS pointed observation}. Since HST was pointed toward the anti-Earth-vector, the image will be especially valuable in searching for slow-moving Kuiper belt objects. It will also be useful for high-redshift galaxy studies. The observations were completed as planned, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The PSEA was configured at 321/1322Z to close the aperture door for any hardware sunpoint entry during the Leonid meteor shower period. The precautions made by HST operations (see yesterday's report and above paragraph for details) for the Leonid meteor shower ensured that the HST passed through the period with no apparent effect on the HST or on operations. Using ROP SR-1, the SSR EDAC error counter was cleared three times. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH